### v1

# On Araki-Type Trace Inequalities

Abstract.
In this paper, we prove a trace inequality $\operatorname{Tr}\left[f(A)A^{s}B^{s}\right]\leq\operatorname{Tr}\big{[}f(A)(A^{1/2}BA^{1/2})^{s}\big{]}$ for any positive and monotone increasing function $f$, $s\in[0,1]$, and positive semi-definite matrices $A$ and $B$. On the other hand, for $s\in[0,1]$ such that the map $x\mapsto x^{s}g(x)$ is positive and decreasing, then $\operatorname{Tr}\big{[}g(A)(A^{1/2}BA^{1/2})^{s}\big{]}\leq\operatorname{Tr}\left[g(A)A^{s}B^{s}\right]$.
1. Introduction
Golden–Thompson trace inequality [1, 2, 3], i.e., for Hermitian matrices $H,K$,
$\operatorname{Tr}\left[\mathrm{e}^{H+K}\right]\leq\operatorname{Tr}\left[\mathrm{e}^{H}\mathrm{e}^{K}\right],$ (1)
and the Araki–Lieb–Thirrling inequality [4, 5]: for $A,B\geq 0$,
$\operatorname{Tr}\left[\left(A^{1/2}BA^{1/2}\right)^{r}\right]\leq\operatorname{Tr}\left[A^{r/2}B^{r}A^{r/2}\right],\quad r\geq 1$ (2)
(the inequality reverses [^1] In some literature, (2) for $r\geq 1$ is called Lieb–Thirrling inequality [4], and the reversed inequality for $r\in[0,1]$ is sometimes called the Araki–Cordes inequality [5, 6]. In this paper, we will use the term, Araki–Lieb–Thirrling inequality, referring to the inequality for all $r\geq 0$. for $r\in[0,1]$) plays important roles in statistical mechanics, quantum information theory, and quantum computing. For example, (1) implies that
$D_{\alpha}(\rho\|\sigma):=\frac{1}{\alpha-1}\log\operatorname{Tr}\left[\rho^{\alpha}\sigma^{1-\alpha}\right]\leq\frac{1}{\alpha-1}\log\operatorname{Tr}\left[\operatorname{\mathrm{e}}^{\alpha\log\rho+(1-\alpha)\log\sigma}\right]=:D_{\alpha}^{\flat}(\rho\|\sigma),\quad\forall\alpha\in(0,1),$ (3)
where $D_{\alpha}(\rho\|\sigma)$ is the Petz–Rényi divergence [7], $D_{\alpha}^{\flat}(\rho\|\sigma)$ is the log-Euclidean Rényi divergence, and $\rho,\sigma$ are density matrices. On the other hand, (2) implies that
$D_{\alpha}^{*}(\rho\|\sigma):=\frac{1}{\alpha-1}\log\operatorname{Tr}\left[\left(\sigma^{\frac{1-\alpha}{2\alpha}}\rho\sigma^{\frac{1-\alpha}{2\alpha}}\right)^{\alpha}\right]\leq D_{\alpha}(\rho\|\sigma),\quad\forall\alpha>0,$ (4)
where $D_{\alpha}^{*}(\rho\|\sigma)$ is the sandwiched Rényi divergence [8, 9]. Essentially, Golden–Thompson inequality and Araki–Lieb–Thirrling inequality manifest the noncommutative nature of quantum mechanics and that certain inequalities instead of the equality will be induced in analytical derivations.
Araki–Lieb–Thirrling trace inequality (2) can be strengthened to the log-majorization relation [5]:
$\left(A^{1/2}BA^{1/2}\right)^{r}\prec_{\log}A^{r/2}B^{r}A^{r/2},\quad r\geq 1$ (5)
(the log-majorization relation reverses for $r\in[0,1]$). Here, for matrices $A,B\in\mathcal{M}_{n}(\mathds{C})$ with nonnegative eigenvalues, the log-majorization of $A$ by $B$, denoted by $A\prec_{\log}B$, means that
$\prod_{i=1}^{k}\lambda_{i}(A)\leq\prod_{i=1}^{k}\lambda_{i}(B),\quad k=1,2\ldots,n,$ (6)
and $\det{A}=\det{B}$, where $\lambda_{i}(A)$ is is $i$-th eigenvalue of $A$ arranged in non-increasing order. Likewise, (1) can be strengthened by (5) and the Lie–Trotter formula [10]:
$\operatorname{\mathrm{e}}^{H+K}\prec_{\log}\left(\operatorname{\mathrm{e}}^{pH/2}\operatorname{\mathrm{e}}^{pK}\operatorname{\mathrm{e}}^{pH/2}\right)^{1/p},\quad\forall p>0.$ (7)
Ando and Hiai later obtained a log-majorization via Löwner–Heinz operator inequality and the antisymmetric tensor power trick (see e.g., [11, Lemma 2.4]), which complements Golden–Thompson type: [10]:
$\left[\mathrm{e}^{pK/2}\left(\mathrm{e}^{-pK/2}\mathrm{e}^{pH}\mathrm{e}^{-pK/2}\right)^{\alpha}\mathrm{e}^{pK/2}\right]^{1/p}\prec_{\log}\operatorname{\mathrm{e}}^{\alpha H+(1-\alpha)K},\quad\alpha\in[0,1],\;p>0.$ (8)
The above then implies that
$D_{\alpha}^{\flat}(\rho\|\sigma)\leq\frac{1}{\alpha-1}\log\operatorname{Tr}\left[\rho\left(\rho^{-1/2}\sigma\rho^{-1/2}\right)^{1-\alpha}\right]=:\widehat{D}_{\alpha}(\rho\|\sigma),\quad\alpha\in(0,1].$ (9)
Bebiano, Lemos and Providência applied Ando and Hiai’s proof technique with Furuta’s operator inequality [12] to obtain the following log-majorization of Araki’s type.
Theorem BLP (BLP Inequality [13]).
For $A,B\geq 0$,
$A^{1+q}B^{q}\prec_{\log}A\left(A^{r/2}B^{r}A^{r/2}\right)^{q/r},\quad 0<q\leq r.$ (10)
Equivalently,
$A^{t+s}B^{s}\prec_{\log}A^{t}\left(A^{1/2}BA^{1/2}\right)^{s},\quad s\in(0,1],\;t>0.$ (11)
Theorem BLP provides an alternative proof to the following trace inequality
$\operatorname{Tr}\left[A\left(\log A+\log B\right)\right]\leq\frac{1}{p}\operatorname{Tr}\left[A\log\left(A^{p/2}B^{p}A^{p/2}\right)\right],\quad p>0$ (12)
(see also the early paper by Hiai and Petz [14], the complementaion by Ando and Hiai [10], and the recent developments by Lieb and Carlen [15]). Inequality (12) then implies that the following relation between Umegaki’s quantum relative entropy $D(\rho\|\sigma)$ [16] and Belavkin–Staszewski’s relative entropy $\widehat{D}(\rho\|\sigma)$ [17], i.e.,
$D(\rho\|\sigma):=\operatorname{Tr}\left[\rho(\log\rho-\log\sigma)\right]\leq\operatorname{Tr}\left[\rho\log\left(\rho^{1/2}\sigma^{-1}\rho^{1/2}\right)\right]=:\widehat{D}(\rho\|\sigma).$ (13)
With efforts to develop the reversed BLP inequalities [18, 19, 20], the above log-majorizations can be unified as the following generalized BLP (GBLP) inequalities.222 Some literature studied different generalizations of the BLP inequality, (11), and used the term generalized BLP [21, 22, 23, 24]. In this paper, we call the inequalities and the reversed versions with the possible range of parameters as in Theorem GBLP as the GBLP inequality.
Theorem GBLP (GBLP Inequality [13, 18, 19, 20]).
For $A,B\geq 0$,
$A^{r+q}B^{q}\prec_{\log}A^{r}\left(A^{p/2}B^{p}A^{p/2}\right)^{q/p},\quad 0<q\leq p,\;r\geq 0.$ (14)
If either $0\leq r\leq p\leq q$ and $p>0$, or $0\leq q<p$ and $-r\geq q$, then (14) holds with reversed log-majorization.
Theorem GBLP and together with a result by Matharu and Aujla (obtained similarly from the Furuta inequality) [25] show the following relations:
$\begin{split}A\left(A^{-1/2}BA^{-1/2}\right)^{s}&\preceq_{\log}A^{1-s}B^{s},\quad s\in[0,1]\cup[2,\infty);\\ A\left(A^{-1/2}BA^{-1/2}\right)^{s}&\succeq_{\log}A^{1-s}B^{s},\quad s\in[1,2].\end{split}$ (15)
The relations given in (15) is intriguing in the following sense. Considering $s\in[0,1]$, then Araki–Lieb–Thirrling inequality shows that $\left(A^{-1/2}BA^{-1/2}\right)^{s}\succeq_{\log}A^{-s}B^{s}$. However, when multiplying both sides by $A$, distributing the power $s$ yields a reversed log-majorization $A\left(A^{-1/2}BA^{-1/2}\right)^{s}\preceq_{\log}A^{1-s}B^{s}$. Hence, the goal of this paper is to identify which function $f$ such that the following trace inequality holds,
$\operatorname{Tr}\left[f(A)A^{s}B^{s}\right]\leq\operatorname{Tr}\left[f(A)\left(A^{1/2}BA^{1/2}\right)^{s}\right],\quad\forall\,s\in[0,1],$ (16)
or when the trace inequality reverses. Obviously, the BLP inequality (11) already provides an answer to (16) for $f(A)=A^{t}$, $t>0$. However, whenever $f$ is not a power function, the core idea of proving Theorems BLP and GBLP, i.e., the antisymmetric tensor power trick with the Furuta inequality, is not applicable.
Our main result is to show that any nonnegative and nondecreasing function $f$ satisfies (16) (see Theorem 4), and, for any $s\in[0,1]$ and any function $g$ such that $x\mapsto x^{s}g(x)$ is nonnegative and nonincreasing, (16) is reversed (see Proposition 5).
We provide our main results and the proofs in Section 2. In Section 3, we discuss the scneario for $s>1$.
2. An Araki-type trace inequality
We first show that (16) holds if $f(A)$ is a kind of orthogonal projection matrices.
Proposition 1.
Let $A=\sum_{i=1}^{n}\lambda_{i}P_{i}$ be the spectral decomposition of positive semi-definite $A$ with $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{n}\geq 0$, and let $\Pi_{k}:=\sum_{i=1}^{k}P_{i}$. Then, for all $B\geq 0$,
[…]
Remark 2.1.
The proposition 1 holds for $\Pi_{k}$ being a projection onto the eigenspace associated with the $k$-th largest eigenvalues. The trace inequality may not be true once replacing $\Pi_{k}$ by an arbitrary eigenprojection $P_{i}$. To see a counter-example, let $s=1/2$ and
$A=\begin{bmatrix}2&0\\ 0&1\end{bmatrix},\,B=\begin{bmatrix}1&1\\ 1&1\end{bmatrix},\,P_{2}=\begin{bmatrix}0&0\\ 0&1\end{bmatrix}.$ (17)
Then, $\operatorname{Tr}[P_{2}A^{s}B^{s}]=1/\sqrt{2}\not\leq\operatorname{Tr}[P_{2}(A^{1/2}BA^{1/2})^{s}]\approx 0.5774$.
In the following proofs and lemmas, we suppose $A>0$, and the case $A\geq 0$ then follows by taking the limit. To prove Proposition 1, we need the following two lemmas.
Lemma 2.
Following the notation in Proposition 1, we denote $I-\Pi_{k}$ by $\Pi_{k}^{\text{c}}$, and let $\tilde{A}_{k}=\Pi_{k}A+\lambda_{k}\Pi_{k}^{\textnormal{c}}=\sum_{i=1}^{n}\max\{\lambda_{i},\lambda_{k}\}P_{i}$. Then, we have
$\operatorname{Tr}\left[\Pi_{k}\left(\tilde{A}_{k}^{1/2}B\tilde{A}_{k}^{1/2}\right)^{s}\right]\leq\operatorname{Tr}\left[\Pi_{k}\left(A^{1/2}BA^{1/2}\right)^{s}\right],\quad\forall\,s\in[0,1].$
The inequality reverses for all $s\in[1,2]$.
Proof.
First, note that $A$ and each $\tilde{A}_{k}$ commute. Let $R=(A\tilde{A}_{k}^{-1})^{1/2}$ be a Hermitian contraction, i.e., $R^{2}\leq I$. Since $x\mapsto x^{s}$ is operator concave for $s\in[0,1]$, the operator Jensen’s inequality [26] implies that
$\left(A\tilde{A}_{k}^{-1}\right)^{1/2}\left(\tilde{A}_{k}^{1/2}B\tilde{A}_{k}^{1/2}\right)^{s}\left(A\tilde{A}_{k}^{-1}\right)^{1/2}=R\left(\tilde{A}_{k}^{1/2}B\tilde{A}_{k}^{1/2}\right)^{s}R\leq\left(R\tilde{A}_{k}^{1/2}B\tilde{A}_{k}^{1/2}R\right)^{s}=\left(A^{1/2}BA^{1/2}\right)^{s}.$
Hence, we obtain
$\operatorname{Tr}\left[\Pi_{k}\left(\tilde{A}_{k}^{1/2}B\tilde{A}_{k}^{1/2}\right)^{s}\right]$ $=\operatorname{Tr}\left[\Pi_{k}\left(A\tilde{A}_{k}^{-1}\right)^{1/2}\left(\tilde{A}_{k}^{1/2}B\tilde{A}_{k}^{1/2}\right)^{s}\left(A\tilde{A}_{k}^{-1}\right)^{1/2}\right]$ (18)
$\leq\operatorname{Tr}\left[\Pi_{k}\left(A^{1/2}BA^{1/2}\right)^{s}\right],$ (19)
with the fact that $\Pi_{k}(A\tilde{A}_{k}^{-1})^{1/2}=(A\tilde{A}_{k}^{-1})^{1/2}\Pi_{k}=\Pi_{k}$.
The case of $s\in[1,2]$ follows similarly by the operator convexity of $x\mapsto x^{s}$, $s\in[1,2]$. ∎
Note that the BLP inequality implies that, for $A,B\geq 0$,
$\operatorname{Tr}\left[A^{t}A^{s}B^{s}\right]\leq\operatorname{Tr}\left[A^{t}\left(A^{1/2}BA^{1/2}\right)^{s}\right],\quad\forall\,s\in[0,1],\,t\geq 0.$ (20)
We have another lemma.
Lemma 3.
For $A>0$ and $B\geq 0$,
$\operatorname{Tr}\left[A^{t}\left(A^{1/2}BA^{1/2}\right)^{s}\right]\leq\operatorname{Tr}\left[A^{t}A^{s}B^{s}\right],\quad\forall\,s\in[0,1],\,t\leq-s.$
Proof.
Substitute $B$ by $A^{-1/2}BA^{-1/2}$ to (20). Then
$\operatorname{Tr}\left[A^{t}A^{s}\left(A^{-1/2}BA^{-1/2}\right)^{s}\right]\leq\operatorname{Tr}\left[A^{t}B^{s}\right],\quad\forall\,s\in[0,1],\,t\geq 0.$
Next, substitute $A$ by $A^{-1}$, then
$\operatorname{Tr}\left[A^{-t-s}\left(A^{1/2}BA^{1/2}\right)^{s}\right]\leq\operatorname{Tr}\left[A^{-t}B^{s}\right],\quad\forall\,s\in[0,1],\;t\geq 0.$
Or,
$\operatorname{Tr}\left[A^{t}\left(A^{1/2}BA^{1/2}\right)^{s}\right]\leq\operatorname{Tr}\left[A^{t}A^{s}B^{s}\right],\quad\forall\,s\in[0,1],\;t\leq-s.$
Proof of Proposition 1.
Combining the two different conditions regarding $t$ in (20) and Lemma 3, we conclude that for $f(x)=x^{t}$, $t\geq 0$ or $f(x)=-x^{t}$, $t\leq-s$, we have
$\operatorname{Tr}\left[f(A)A^{s}B^{s}\right]\leq\operatorname{Tr}\left[f(A)\left(A^{1/2}BA^{1/2}\right)^{s}\right],\quad\forall\,s\in[0,1].$ (21)
Notice that $\operatorname{Tr}[\Pi_{k}A^{s}B^{s}]=\operatorname{Tr}[\Pi_{k}\tilde{A}_{k}^{s}B^{s}]$ for $\Pi_{k}$ introduced in Lemma 2. Hence, to complete the proof of Proposition 1, by Lemma 2 we only need to show that
$\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}^{s}B^{s}\right]\leq\operatorname{Tr}\left[\Pi_{k}\left(\tilde{A}_{k}^{1/2}B\tilde{A}_{k}^{1/2}\right)^{s}\right],\quad\forall\,s\in[0,1].$ (22)
To show this, we first introduce $\tilde{A}_{k,\varepsilon}=\Pi_{k}A+(\lambda_{k}-\varepsilon)\Pi_{k}^{\mathrm{c}}$ as an approximation of $\tilde{A}_{k}$, for $\varepsilon\in(0,\lambda_{k})$, where $\lambda_{k}$ is $k$-th largest eigenvalue of $A$.
Define
$g_{t}(x)=1-\left(\frac{x}{\lambda_{k}-\varepsilon}\right)^{t},\quad\forall\,t\leq-s,$
then, (21) implies that
$\operatorname{Tr}\left[g_{t}(\tilde{A}_{k,\varepsilon})\tilde{A}_{k,\varepsilon}^{s}B^{s}\right]\leq\operatorname{Tr}\left[g_{t}(\tilde{A}_{k,\varepsilon})\left(\tilde{A}_{k,\varepsilon}^{1/2}B\tilde{A}_{k,\varepsilon}^{1/2}\right)^{s}\right],\quad\forall\,s\in[0,1],\,t\leq-s,$
with noticing that the constant function $1$ is just $f(x)=x^{t}$ at $t=0$.
Let $t\to-\infty$, then $g_{t}(\tilde{A}_{k,\varepsilon})\to\Pi_{k}$, and hence
$\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k,\varepsilon}^{s}B^{s}\right]\leq\operatorname{Tr}\left[\Pi_{k}\left(\tilde{A}_{k,\varepsilon}^{1/2}B\tilde{A}_{k,\varepsilon}^{1/2}\right)^{s}\right],\quad\forall\,s\in[0,1].$
Finally, by letting $\varepsilon\to 0$, we obtain
$\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}^{s}B^{s}\right]\leq\operatorname{Tr}\left[\Pi_{k}\left(\tilde{A}_{k}^{1/2}B\tilde{A}_{k}^{1/2}\right)^{s}\right],\quad\forall\,s\in[0,1].$
In conclusion, for all $s\in[0,1]$, we have
$\operatorname{Tr}\left[\Pi_{k}A^{s}B^{s}\right]=\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}^{s}B^{s}\right]\leq\operatorname{Tr}\left[\Pi_{k}\left(\tilde{A}_{k}^{1/2}B\tilde{A}_{k}^{1/2}\right)^{s}\right]\leq\operatorname{Tr}\left[\Pi_{k}\left(A^{1/2}BA^{1/2}\right)^{s}\right],$
which completes the proof of Proposition 1. ∎
Now we are in a position to prove our main results.
Theorem 4.
Let $A,B\geq 0$. For any nonnegative and nondecreasing function $f$ on an interval $\mathcal{J}$, where $\operatorname{spec}(A)\subseteq\mathcal{J}$, we have
$\operatorname{Tr}\left[f(A)A^{s}B^{s}\right]\leq\operatorname{Tr}\left[f(A)\left(A^{1/2}BA^{1/2}\right)^{s}\right],\quad\forall\,s\in[0,1].$
Proof.
Since $f$ is nonnegative and nondecreasing, we can choose some appropriate nonnegative coefficients $c_{k}$ for each $k=1,\cdots,n$, such that $f(A)=\sum_{k=1}^{n}c_{k}\Pi_{k}$. Then, the claim follows from Proposition 1. ∎
As above, we may also apply substitutions to obtain another direction of the same inequality, but with a different condition.
Proposition 5.
Let $A,B\geq 0$. For any $s\in[0,1]$ and any function $g$ on an interval $\mathcal{J}$, where $\operatorname{spec}(A)\subseteq\mathcal{J}$, such that $x\mapsto x^{s}g(x)$ is nonnegative and nonincreasing, then
$\operatorname{Tr}\left[g(A)\left(A^{1/2}BA^{1/2}\right)^{s}\right]\leq\operatorname{Tr}\left[g(A)A^{s}B^{s}\right].$
Proof.
Again, we assume $A>0$, and extend to $A\geq 0$ via a limit. Set $f(x)=x^{-s}g(x^{-1})$, which is nonnegative and nondecreasing. Since the domain of $f$ includes the spectrum of $A^{-1}$, by Theorem 4,
$\operatorname{Tr}\left[g(A)\left(A^{{1}/{2}}BA^{{1}/{2}}\right)^{s}\right]$ $=\operatorname{Tr}\left[f\left(A^{-1}\right)\left(A^{-1}\right)^{s}\left(A^{{1}/{2}}BA^{{1}/{2}}\right)^{s}\right]$ (23)
$\leq\operatorname{Tr}\left[f\left(A^{-1}\right)\left(A^{-1/2}\cdot A^{{1}/{2}}BA^{{1}/{2}}\cdot A^{-1/2}\right)^{s}\right]$ (24)
$=\operatorname{Tr}\left[g(A)A^{s}B^{s}\right].$ (25)
3. A Conjecture for $s>1$
Theorem 4 shows the trace inequality for all $s\in[0,1]$. Inspired by the reverse BLP inequality [18, 21, 22], we conjecture the following reverse direction.
Conjecture 1.
For any nonnegative and nondecreasing function $f$ on an interval $\mathcal{J}$, where $\operatorname{spec}(A)\subseteq\mathcal{J}$, we have
$\operatorname{Tr}\left[f(A)A^{s}B^{s}\right]\geq\operatorname{Tr}\left[f(A)\left(A^{1/2}BA^{1/2}\right)^{s}\right],\quad\forall\,s>1.$
Note that $x\mapsto x^{s}$ is operator convex for $s\in[1,2]$, we may adapt the previous technique to prove Conjecture 1 for $s\in[1,2]$, provided that the condition $0\leq r\leq p\leq q$ in the reverse direction of Theorem GBLP can be relaxed to $0\leq p\leq q$. Thereby, we conjecture the following reverse log-majorization, which complements (14).
Conjecture 2.
For $A,B\geq 0$,
$\operatorname{Tr}\left[A^{r+q}B^{q}\right]\geq\operatorname{Tr}\left[A^{r}\left(A^{p/2}B^{p}A^{p/2}\right)^{q/p}\right],\quad 0\leq p\leq q,\;r\geq 0.$ (26)
(To show Conjecture 1 for $s\in[1,2]$, the trace inequality version of Conjecture 2 would be sufficient.)
In fact, we can directly prove Conjecture 1 for a special case of $s=2$ without resorting to Conjecture 2.
Proposition 6.
For any nonnegative and nondecreasing function $f$ on an interval $\mathcal{J}$, where $\operatorname{spec}(A)\subseteq\mathcal{J}$, we have
$\operatorname{Tr}\left[f(A)A^{s}B^{s}\right]\geq\operatorname{Tr}\left[f(A)\left(A^{1/2}BA^{1/2}\right)^{s}\right],\quad s=2.$
Proof.
Our claim is based on Lemma 2 and the reverse direction to (22), which is presented in the following Lemma 7. ∎
Lemma 7.
Let $A=\sum_{i=1}^{n}\lambda_{i}P_{i}$ be the spectral decomposition of positive semi-definite $A$ with $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{d}\geq 0$, $\Pi_{k}:=\sum_{i=1}^{k}P_{i}$, $\Pi_{k}^{\textnormal{c}}:=I-\Pi_{k}$, and $\tilde{A}_{k}:=\Pi_{k}A+\lambda_{k}\Pi_{k}^{\textnormal{c}}$. Then,
$\operatorname{Tr}\left[\Pi_{k}\left(\tilde{A}_{k}^{1/2}B\tilde{A}_{k}^{1/2}\right)^{s}\right]\leq\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}^{s}B^{s}\right],\quad s=2.$ (27)
Proof.
Using the fact that $\lambda_{k}\Pi_{k}\leq\Pi_{k}\tilde{A}_{k}\Pi_{k}$ by construction and the Araki–Lieb–Thirrling trace inequality (2), we calculate
$\operatorname{Tr}\left[\Pi_{k}\left(\tilde{A}_{k}^{1/2}B\tilde{A}_{k}^{1/2}\right)^{2}\right]$ $=\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}\Pi_{k}B{\tilde{A}_{k}}B\right]$ (28)
$=\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}\Pi_{k}B{\left(\Pi_{k}\tilde{A}_{k}\Pi_{k}+\lambda_{k}\Pi_{k}^{\text{c}}\right)}B\right]$ (29)
$=\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}\Pi_{k}B{\Pi_{k}\tilde{A}_{k}\Pi_{k}}B\right]+\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}\Pi_{k}B{(\lambda_{k}\Pi_{k}^{\text{c}})}B\right]$ (30)
$=\operatorname{Tr}\left[(\Pi_{k}\tilde{A}_{k}\Pi_{k}\Pi_{k}B\Pi_{k})^{2}\right]+\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}\Pi_{k}B{(\lambda_{k}\Pi_{k}^{\text{c}})}B\right]$ (31)
$=\operatorname{Tr}\left[(\Pi_{k}\tilde{A}_{k}\Pi_{k}\Pi_{k}B\Pi_{k})^{2}\right]+\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}\Pi_{k}{\lambda_{k}\Pi_{k}}B{\Pi_{k}^{\text{c}}}B\right]$ (32)
$\leq\operatorname{Tr}\left[(\Pi_{k}\tilde{A}_{k}\Pi_{k}\Pi_{k}B\Pi_{k})^{2}\right]+\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}\Pi_{k}{\Pi_{k}\tilde{A}_{k}\Pi_{k}}B{\Pi_{k}^{\text{c}}}B\right]$ (33)
$=\operatorname{Tr}\left[(\Pi_{k}\tilde{A}_{k}\Pi_{k}\Pi_{k}B\Pi_{k})^{2}\right]+\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}^{2}\Pi_{k}B{\Pi_{k}^{\text{c}}}B\right]$ (34)
(2)${\leq}\operatorname{Tr}\left[(\Pi_{k}\tilde{A}_{k}\Pi_{k})^{2}(\Pi_{k}B\Pi_{k})^{2}\right]+\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}^{2}\Pi_{k}B{\Pi_{k}^{\text{c}}}B\right]$ (35)
$=\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}^{2}\Pi_{k}B\Pi_{k}B\right]+\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}^{2}\Pi_{k}B{\Pi_{k}^{\text{c}}}B\right]$ (36)
$=\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}^{2}\Pi_{k}B(\Pi_{k}+{\Pi_{k}^{\text{c}}})B\right]$ (37)
$=\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}^{2}B^{2}\right].$ (38)
4. Conclusions
In this paper, we derived some Araki-type trace inequalities that involve monotone functions other than powers. Although usual Araki-type inequalities and the BLP inequalities are expressed in a stronger form of log-majorization, however, the antisymmetric tensor power trick may not be applicable for functions beyond powers. On the other hand, trace inequalities find more and immediate applications in other fields such as quantum information science and mathematical physics, which motivates us to further Araki-type trace inequalities to a broader form.
Numerical experiments suggest that Conjecture 1 for $s>1$ and a slightly general reverse BLP inequality in Conjecture 2 could be true, which is left for future work.